### v1

@@ Abstract
Image augmentation constitutes a common and convenient way to mitigate issues stemming from dataset limitations, by automatically creating additional variants of each training image and utilizing them for enhancing the training set [172]. Typically, the generated variants exhibit differences in appearance, but retain semantic content identical to that of the original image. Extending the training dataset with such synthetic images increases its diversity and improves, in many cases, the learning and recognition performance of DNNs that are being trained on it [205]. This behaviour stems from image augmentation essentially acting as an additional regularizing mechanism while training the DNN and, as a result, helping to prevent overfitting [108, 136].
Traditional approaches for image augmentation, such as geometric transformations (e.g., image rotation, flip, crop, scaling, horizontal/vertical translation, squeezing, etc.) and color space adjustments or photometric transforms (e.g., blurring, sharpening, jittering, etc.) are still very common [173, 136, 179]. Multiple transformations of this type can be composed together, so that an even wider set of augmented images can be generated from the original dataset. These methods leverage domain knowledge to produce synthetic examples similar to the initial ones. More recently proposed image augmentation methods in this general vein are a set of strategies for systematically corrupting the original images, in order to generate augmented variants. This category of methods includes, among others: a) ‘mixup’ [192], which uses convex combinations of pairs of training images and their labels, b) ‘cutout’ [37], which randomly masks square regions of an input image, c) ‘cutmix’ [186], which randomly combines two training images by masking the first with a region of the second (and vice versa), and d) ‘patchshuffle’ [70], which uses a kernel filter to randomly swap the pixel values in a sliding window.
@@
Unlike traditional methods, which manipulate existing images to generate variants, Diffusion Models (DMs) can be readily exploited for image augmentation practices, by synthesizing new, realistically-looking and plausible images [199, 174, 5]. DMs constitute a sophisticated class of generative DNNs that excel in implicitly modeling the underlying data generating distribution and the structure of complex images. This capability allows them to essentially sample fake novel images from their training dataset’s distribution, which are simultaneously diverse, highly realistic and representative of unseen data scenarios, as they encompass subtle details and preserve the inherent structure of the original dataset [196, 149]. Thus, they can be directly utilized for meaningfully augmenting the latter.
@@
Despite the recent progress and achievement on the application of DMs for advanced image augmentation, there are very few relevant surveys in the literature. Existing surveys either focus on traditional image augmentation [179] or provide a general overview of DMs without delving into their specific application for image augmentation [18, 177, 141]. For instance, the study of [99] presents a comprehensive taxonomy of image augmentation approaches, including input space transformations, feature space augmentation, data synthesis and meta-learning based methods. However, it does not cover the latest advancements concerning the use of DMs. In contrast, the work of [35] presents three generic diffusion modeling frameworks, which are based on denoising diffusion probabilistic models, noise conditioned score networks, and stochastic differential equations, but does not explore their potential for image augmentation. Some surveys have touched upon the efficiency aspect of DMs [151] or their application in specific domains like medical imaging [72, 73]. However, these works do not provide a comprehensive overview of DMs for augmentation across various computer vision tasks. Furthermore, a recent survey categorizes augmentation methods based on large learning models, including those based on DMs, but does not focus specifically on them [206].
@@ 2.4 Diffusion Models in Latent Space
2.5 Foundation Diffusion Models for Image Synthesis
Conditional LDMs have boosted the development of the ‘Stable Diffusion’ (SD) Foundation Model, a Text-to-Image generator (T2I) which was pretrained on the LAION-5B dataset [132]. SD has dominated much of recent research related to generative image synthesis for image augmentation. Still, several attempts have been made to further improve it. For instance, ‘Stable Diffusion XL’ (SDXL) [110] innovates over the basic SD in three ways:
It boasts a U-Net three times more complex and leverages a dual text encoding system for text conditioning. This new text encoder (OpenCLIP ViT-bigG/14), operating alongside the original one, significantly expands the model’s capacity.
It enhances control over the final image crop, via the incorporation of size- and crop-conditioning during training. This is implemented by feeding crop parameters to the model as conditioning parameters via Fourier feature embeddings.
Its inference stage operates in two steps: a ‘base’ model generates an initial image, which is then fed to a ‘refiner’ model that adds finer, higher-quality details.
Another advanced variant is PixArt- [25], which differentiates from SD in three aspects:
Training strategy decomposition: Three distinct training steps are devised to respectively optimize pixel dependency, text-image alignment, and image aesthetic quality.
Efficient T2I Transformer: Cross-attention modules are incorporated into the Diffusion Transformer (DiT) to inject text conditions and streamline the computation-intensive class-conditioning branch.
Highly informative data: Emphasizing the significance of concept density in text-image pairs, a large Vision-Language Model (VLM) is leveraged to auto-label dense pseudo-captions, assisting in text-image alignment learning.
More recently, the authors of [43] have proposed ‘Stable Diffusion 3’ (SD3), which relies on a Transformer neural architecture [107], instead of a U-Net Convolutional Neural Network (CNN). It utilizes a distinct set of parameters for the two involved modalities, namely text and images; thus, allowing for a two-way exchange of information between images and text tokens, making use of an attention mechanism.
@@ 3 Taxonomy of Diffusion Models for Image Augmentation
Complementarily to the categorization described above, Fig. 3 illustrates a timeline representation that contains key recent DM-powered image augmentation methods. Each entry corresponds to a critical milestone work that significantly impacted the research field. Naturally, more recent works are associated with more complex and advanced DM models/architectures, leading also to superior performance. Moreover, Table 3 compactly summarizes the most important works, taking into account the category and subcategory to which they belong.
@@ 4.1.4 I2I (Image-to-Image) Translation
Several methods make use of conditional DMs for I2I translation. For example, SDEdit [96] utilizes a conditional DM, guided by stochastic differential equations, to perform image editing. CycleNet [174] and DiffusionCLIP [74] finetune a conditional DM with a CLIP-based loss to ensure that the generated image matches the target image’s text description. Palette et al. [123] employ conditional DMs to learn the distribution for various tasks, like colorization, inpainting, JPEG restoration, and uncropping. The model incorporates a U-Net architecture with self-attention layers, conditioning on the input image through concatenation. During training, it predicts the noise added to the original image, minimizing L2 or L1 loss between the predicted and the actual noise. Then, the inference process involves iterative denoising over 1000 timesteps, starting from Gaussian noise and progressively refining the output. The method’s strength lies in its ability to handle multiple tasks with a single architecture, eliminating the need for task-specific customizations (Fig. 7).
A different group of methods focus on disentangling style and content for more targeted I2I translation. For instance, the approach of [82] makes use of separate encoders for style and content, and then injects these representations into the DM to generate the translated image. Other methods explore few-shot or zero-shot I2I translation. For example, pix2pix-zero [106] first inverts the input image to obtain a noise map, using DDIM and BLIP. These noise maps are then regularized to improve editability, using an autocorrelation objective. Edit directions are automatically discovered by generating diverse sentences for the source and target domains, and computing the mean difference between their CLIP embeddings. In order to preserve image structure during editing, a novel cross-attention guidance technique is employed, which matches the edited cross-attention maps with reference maps corresponding to the original structure. The method builds upon the SD model, utilizing its CLIP conditioning and cross-attention layers. On the other hand, Dual Diffusion Implicit Bridges (DDIBs) [145] encompass two independently trained DMs (one for the source domain and one for the target one) to perform I2I translation, without requiring joint training on paired data.
@@ 4.1.5 Counterfactual Augmentation
A great portion of methods focus on generating counterfactuals for specific applications. For example, MEDJOURNEY [52] generates counterfactual medical images by conditioning the DM on a text description of the desired change, while the methods of [95] and [185] generate counterfactual images to explain the model’s behavior and to improve its robustness. On the other hand, a different set of approaches utilize counterfactual augmentation to address bias and fairness issues in datasets and models. For example, the work of [105] utilizes a conditional DM to generate counterfactual images that balance the distribution of sensitive attributes (e.g., gender, race) in the dataset. The model is conditioned on the desired attribute distribution and generates images that match it, while preserving the remaining aspects of the image.
@@ 4.3.2 Other Domain-Specific Applications
Object Detection: The approaches of [44] and [194] leverage DMs to generate synthetic data, significantly enhancing model training and performance. These methods make use of conditional DMs trained on object detection datasets and generate new samples, by conditioning on the desired object categories and bounding boxes.

### v2

@@ Abstract
Image augmentation constitutes a common preprocessing step in machine learning workflows, designed to increase the visual diversity of the training dataset, without introducing additional independent images. The goal is to enhance the learning process, while operating within a fixed set of data [189]. Augmentation addresses limitations in dataset size, by automatically creating additional variants of existing training images. These variants are generated by directly modifying the original images, ensuring that the transformed images differ in appearance, but retain their semantic content [234]. In contrast, general image generation (or image synthesis) involves sampling from a model that approximates the overall data distribution of the dataset, rather than transforming individual training images. This approach generates entirely new images that are not directly derived from specific originals, but they are instead representative of the underlying characteristics of the whole dataset [22]. Extending the dataset with such augmented images, either synthesized or modified, increases its diversity and improves, in many cases, the downstream learning and recognition performance of DNNs that are being trained using it [272]. This behaviour stems from image augmentation essentially acting as an additional regularizing mechanism while training the DNN and, as a result, helping to prevent overfitting [152, 189].
Traditional approaches for image augmentation, such as geometric transformations (e.g., image rotation, flip, crop, scaling, horizontal/vertical translation, squeezing, etc.) and color space adjustments or photometric transforms (e.g., blurring, sharpening, jittering, etc.) are still very common [235, 189, 242]. Multiple transformations of this type can be composed together, so that an even wider set of augmented images can be generated from the original dataset. These methods leverage domain knowledge to produce synthetic examples similar to the initial ones. More recently proposed image augmentation methods in this general vein are a set of strategies for systematically corrupting the original images, in order to generate augmented variants. This category of methods includes, among others: a) ‘Mixup’ [258], which uses convex combinations of pairs of training images and their labels, b) ‘Cutout’ [49], which randomly masks square regions of an input image, c) ‘Cutmix’ [250], which randomly combines two training images by masking the first with a region of the second (and vice versa), d) ‘Patchshuffle’ [96], which uses a kernel filter to randomly swap the pixel values in a sliding window, e) ‘Copy-Paste’ [58], which pastes segmented object instances onto random background images, f) ‘Co-Mixup’ [102], which optimizes saliency-guided mixing of multiple inputs while encouraging diversity among outputs, g) ‘RandAugment’ [46], which randomly selects transformations from a predefined set, each applied with magnitude , eliminating the need for separate policy search, h) ‘GuidedMixup’ [97], which uses saliency maps to guide image mixing while preserving salient regions through pixel-wise mixing ratios, and i) ‘CAL-AUG’ [164], which randomly replaces learned attention maps with counterfactual ones.
@@
Unlike traditional methods, which directly manipulate existing images to generate variants, Diffusion Models (DMs) can be exploited for image augmentation, by learning to synthesize new, realistically-looking and plausible images, given a training dataset [265, 236, 6]. These are general image generation algorithms, aiming to learn a model of the underlying distribution of the selected training dataset and then allow its sampling, in order to synthesize novel images. These generated images do not have a direct one-to-one correspondence with the original training ones, but instead reflect the dataset’s underlying statistical patterns. Generative methods can be used for various tasks, such as procedural content generation and simulation, but they can also implement dataset expansion, as a specific type of image augmentation at the dataset level. Additionally, when coupled with a conditioning mechanism, they can replace traditional augmentation approaches and synthesize modified variants of specific given images, through conditional generation that allows one-to-one correspondence between an original training image and a synthesized variant. DMs are a sophisticated class of generative DNNs that excel in implicitly modeling the underlying data generating distribution and the structure of complex images. This capability allows them to essentially sample fake novel images from their training dataset’s distribution, which are simultaneously diverse, highly realistic and representative of unseen data scenarios, as they encompass subtle details and preserve the inherent structure of the original dataset [262, 207]. Thus, they can be utilized for meaningfully augmenting the latter.
@@
Despite the increasing importance of DMs in image augmentation, there is a gap in the relevant literature, since no existing review focuses specifically on this application and its particularities. Published surveys on DMs do not focus on image augmentation in a targeted manner, given the recent progress and achievements in the field. Existing surveys either focus on traditional image augmentation [242], or provide a general overview of DMs without delving into their specific application for image augmentation [22, 240, 195]. For instance, the study of [140] presents a comprehensive taxonomy of image augmentation approaches, including input space transformations, feature space augmentation, data synthesis and meta-learning based methods. However, it does not cover the latest advancements concerning the use of DMs. In contrast, the work of [45] presents three generic diffusion modeling frameworks, which are based on denoising diffusion probabilistic models, noise conditioned score networks, and stochastic differential equations, but does not explore their potential for image augmentation. Some surveys have touched upon the efficiency aspect of DMs [209] or their application in specific domains like medical imaging [99, 100]. However, these works do not provide a comprehensive overview of DMs for augmentation across various computer vision tasks. Furthermore, a recent survey categorizes augmentation methods based on large learning models, including those based on DMs, but does not focus specifically on them [273].
The above-described situation has motivated this study, in an attempt to remedy the identified gap. Its main contribution is a systematic, comprehensive and in-depth review of DM-based approaches specifically for image augmentation, covering a wide range of strategies, tasks and applications. In particular, the current work analyzes fundamental principles, introduces a detailed taxonomy of methods, examines evaluation metrics, discusses current challenges and provides perspectives on future research directions.
@@ 2.4 Diffusion Models in Latent Space
2.5 Types of Conditioning
DMs can be conditioned on various types of information to guide the generation process. The most common types of conditioning include:
Class Label Conditioning: Models can be conditioned on discrete class labels [50], enabling class-specific image generation. This is often implemented through class embeddings that are concatenated with or added to the model’s intermediate features.
Segmentation Map Conditioning: Incorporating segmentation maps allows for control over the spatial layout and the structure of the generated images. Methods like [218] incorporate segmentation maps through specialized encoders of feature injection to control the spatial distribution of semantic classes. This enables precise control over object placement and scene composition.
Pose and Structure Conditioning: Human pose maps [259], edge maps, or semantic segmentation masks can guide the spatial structure of the generated images. These are typically processed through specialized encoders and injected into the model via cross-attention or feature concatenation.
The most suitable type of conditioning depends on the specific application and the desired degree of control over the generation procedure. Recent work has focused on developing more sophisticated approaches to conditioning that enable finer-grained control, while maintaining generation quality [259, 269].
2.6 Diffusion Transformer
The Diffusion Transformer (DiT) [151] architecture represents a significant departure from U-Net-based designs, replacing convolutional neural layers with Transformer blocks for processing visual data. DiT maintains the same high-level structure as previous DMs, but processes images as sequences of spatial patches in the vein of Vision Transformers (ViT) [53]. The architecture consists of: a) A patchification layer that splits images into non-overlapping patches and linearly projects them, b) A sequence of Transformer blocks with self-attention mechanisms, c) Time step and condition embeddings that are added to the patch embeddings, and d) A final decoder layer that reconstructs the spatial input image. The key advantages of DiT over U-Net-LDM include better scaling properties, improved handling of long-range dependencies through self-attention, and more flexible integration of conditioning information [75, 137]. However, DiT typically requires more computational resources for training, due to the quadratic complexity of the self-attention operations. Recent work has shown that DiT-based models can achieve superior generation quality when scaled to sufficient size, leading to their wide adoption [151, 62].
2.7 Foundation Diffusion Models for Image Generation
Model | Release Date | License | Conditioning Mechanisms | Text Encoder | Training Data | Params | Res. (pixels) | Notable Features | Indicative Derived Methods |
Pixel-Space U-Net FDMs | |||||||||
Imagen [173] | May’22 | Prop. | Text, Img | T5-XXL [159] | LAION-400M + Prop. ( 860M) | 2B | First that explored using frozen LLMs as text encoders | [217, 34, 84] | |
DeepFloyd IF [198] | Apr’24 | Open | Text, Img | T5-XXL [159] | LAION-5B (1B) [183] | 400M - 4.3B | Cascaded pixel diffusion | [202] | |
Latent-Space U-Net FDMs | |||||||||
[165] | Dec’21 | Open | Text, Img, Mask, Layout, Semantic Map | [158] | LAION-5B [183] | First major open-source LDM | [156, 23, 224] | ||
[161] | Apr’22 | Prop. | Text, CLIP Embeddings | [158] | Prop. (M) | Two-stage generation using CLIP latents | [14] | ||
[165] | Nov’22 | Open | Text, Img, Mask, Layout, Semantic Map | [158] | LAION-5B [183] | 860M | Improved text understanding | [259, 257] | |
SD Cascade [153] | Feb’23 | Open | Text, Img, Mask, Layout, Semantic Map, Edge Map | CLIP-H [88] | LAION-5B [183] (103M) | 4B | Highly compressed semantic latent space | [154] | |
SDXL [156] | Jul’23 | Open | Text, Img | CLIP ViT-L [158] + OpenCLIP ViT-bigG [39] | Prop. | 2.6B | Dual text encoders with size and crop conditioning | [180] | |
Kandinsky 3 [5] | Oct’23 | Open | Text, Img, Mask | Flan-UL2 [205] | LAION-5B + COYO-700M + Prop. (150M) | 11.9B | Single-stage pipeline with large text encoder for better text understanding | [150] | |
[14] | Oct’23 | Prop. | Text, Img | [159] | Prop. (B) | - | Training on AI-generated detailed captions improves prompt following ability | - | |
SDXL Turbo [180] | Feb’24 | Open | Text, Img | CLIP-ViT-g-14 [158] | Distilling from SDXL | 3.5B | Single-step generation | [199] | |
Transformer-Based FDMs | |||||||||
PixArt- [29] | Sep’23 | Open | Text | Flan-T5-XXL [159] | SAM-LLAVA + JourneyDB + Prop. (25M) | 600M | Efficient training with only 1% cost of SOTA models | [30] | |
PixArt- [30] | Jan’24 | Open | Text, Img | T5 [159] | Prop (120K) | 600M | Fast 4-step sampling | - | |
SD 3 [59] | Mar’24 | Mixed | Text, Img | T5 XXL [159] | Prop. (B) | Rectified flow formulation enables efficient few-step sampling | - |
The evolution of DMs for image generation has been marked by several key architectural innovations and training strategies. These developments have resulted in the proliferation of commonly used Foundation Models (FMs) for image synthesis, which are available in a pretrained form and can be employed as a basis for more specialized methods. The architectures can be primarily categorized into U-Net-based [165] and Transformer-based [151] ones, with U-Net-based methods further divided into those operating directly in pixel space versus those performing inference in latent spaces. All methods operate in the latent space of a pretrained autoencoder, unless those that perform directly in the pixel space. Moreover, U-Net-based methods employ a conditional LDM as backbone (or DDPM [81], if they operate in pixel space), while Transformer-based approaches utilize a conditional DiT as their backbone.
Table 1 illustrates key Foundation Diffusion Models (FDMs) presented in the literature, emphasizing on the following important characteristics: a) Release date, b) License, c) Conditioning mechanisms, d) Text encoder, e) Training dataset, f) Number of parameters, g) Resolution (in pixels), h) Notable features, and i) Indicative/important methods relying on them. As can be seen from Table 1, most FDMs operate in latent space and use text/image conditioning, while U-Net is the dominant neural architecture before the recent emergence of Transformer-based approaches. LAION-5B [183] and its variants serve as the predominant training dataset for many open-source FDMs, highlighting its significance in image generation. There is also an evolution in text encoders from CLIP-based models towards more sophisticated language models like T5-XXL, suggesting a trend toward better text understanding.
@@ 3 Taxonomy of Diffusion Models for Image Augmentation
Complementarily to the categorization described above, Fig. 3 illustrates a timeline representation that contains key recent DM-powered image augmentation methods. Each entry corresponds to a critical milestone work that significantly impacted the research field. Naturally, more recent works are associated with more complex and advanced DM models/architectures, leading also to superior performance. Moreover, Table 3 illustrates key representative methods, belonging to each category defined in the taxonomy of Fig. 2.
@@ 4.1.4 I2I (Image-to-Image) Translation
Several methods make use of conditional DMs for I2I translation [136, 160]. For example, SDEdit [135] utilizes a conditional DM, guided by stochastic differential equations, to perform image editing. CycleNet [236] and DiffusionCLIP [101] finetune a conditional DM with a CLIP-based loss to ensure that the generated image matches the target image’s text description. Palette et al. [172] employ conditional DMs to learn the distribution for various tasks, like colorization, inpainting, JPEG restoration, and uncropping. The model incorporates a U-Net architecture with self-attention layers, conditioning on the input image through concatenation. During training, it predicts the noise added to the original image, minimizing L2 or L1 loss between the predicted and the actual noise. Then, the inference process involves iterative denoising over 1000 timesteps, starting from Gaussian noise and progressively refining the output. The method’s strength lies in its ability to handle multiple tasks with a single architecture, eliminating the need for task-specific customizations (Fig. 7).
A different group of methods focus on disentangling style and content for more targeted I2I translation [113, 124]. For instance, the approach of [113] makes use of separate encoders for style and content, and then injects these representations into the DM to generate the translated image. Other methods explore few-shot or zero-shot I2I translation. For example, pix2pix-zero [149] first inverts the input image to obtain a noise map, using DDIM and BLIP. These noise maps are then regularized to improve editability, using an autocorrelation objective. Edit directions are automatically discovered by generating diverse sentences for the source and target domains, and computing the mean difference between their CLIP embeddings. In order to preserve image structure during editing, a novel cross-attention guidance technique is employed, which matches the edited cross-attention maps with reference maps corresponding to the original structure. The method builds upon the SD model, utilizing its CLIP conditioning and cross-attention layers. On the other hand, Dual Diffusion Implicit Bridges (DDIBs) [200] encompass two independently trained DMs (one for the source domain and one for the target one) to perform I2I translation, without requiring joint training on paired data.
@@ 4.1.5 Counterfactual Augmentation
A great portion of methods focus on generating counterfactuals for specific applications. For example, the methods of [133] and [249] generate counterfactual images to explain the model’s behavior and to improve its robustness. On the other hand, a different set of approaches utilize counterfactual augmentation to address bias and fairness issues in datasets and models. For example, the work of [147] utilizes a conditional DM to generate counterfactual images that balance the distribution of sensitive attributes (e.g., gender, race) in the dataset. The model is conditioned on the desired attribute distribution and generates images that match it, while preserving the remaining aspects of the image.
@@ 4.2.2 Adaptation Methods
The methods of [89] and [90] introduce a prompt-guided approach that enhances both in- and cross-domain image classification. Specifically, they leverage image editing with carefully designed conditional prompts to generate augmented images based on concatenation of original and edited content, followed by blending with fractal patterns, which reduces overfitting and ensures robust training.
@@ 4.3.2 Other Domain-Specific Applications
Object Detection: The approaches of [61], [260], [120], [204] and [130] leverage DMs to generate synthetic data, significantly enhancing model training and performance. These methods make use of conditional DMs trained on object detection datasets and generate new samples, by conditioning on the desired object categories and bounding boxes.
4.4 Comparison of DM-Based Methods for Image Augmentation
Having presented the various methodologies for DM-based image augmentation, this section realizes a detailed and in-depth comparison among the most important methods of each category, in order to provide further critical insights. In particular, Table 3 analyzes several key/important methods, focusing on the following aspects: a) The particular task to be performed, b) If/which FDM is used, c) Input requirements, d) Key exhibited innovation, e) Notable limitations, and f) Fine-tuning requirements. The considered methods, which cover all main categories of the proposed taxonomy (presented in Fig. 2 and Table 3) were selected based on the following criteria: a) High citation impact and community adoption, b) Introduction of novel technical approaches or architectures, and c) Demonstration of significant practical capabilities or limitations that influenced subsequent research works.
Method | Task | FDM | Input Requirements | Key Innovation | Notable Limitations | Fine-tuning Requirements |
Semantic Manipulation | ||||||
[238] | I2I Translation | SD | Source image + reference image + mask | Content bottleneck prevents trivial copying | May not preserve fine details | Self-supervised training on OpenImages |
[101] | I2I Translation | - | Image + text prompt | Multi-attribute control via noise combination | Limited to pretrained domains | Reference images for domain finetuning |
[127] | Inpainting | - | Image + binary mask | Resampling strategy for mask-agnostic inpainting | Significantly slower than GAN-based methods | - |
[135] | Editing | - | User strokes or patches | No task-specific training needed | Quality depends on initial noise level | - |
[98] | Editing | Imagen | SD | Image + text prompt | Linear semantic embedding interpolation | Slow optimization (8min/image) | Target text embedding optimization |
[21] | Instruction-based Editing | SD v1.5 | Image + text instruction | Direct instruction followed without optimization | Cannot handle spatial reasoning | Image-instruction-output triplets training |
[8] | Local Editing | - | Image + mask + text prompt | Spatially blended noise guidance for local editing | 30s generation time per image | - |
[37] | Object Teleportation | SD 2.1 | Target object + scene image + location box | Zero-shot object teleportation with identity preservation | Small detail preservation issues | Light UNet decoder fine-tuning |
[85] | Multi-condition Generation | - | Multiple optional conditions (text/sketch/depth/etc.) | Composable conditions with exponential control space | Performance tradeoff in single-condition scenarios | 60M image multi-task training |
[259] | Image Control | SD 1.5/2.1 | Image + text prompt | Zero convolution prevents catastrophic forgetting | Higher GPU memory and training time | Trainable copy with zero convolution |
Personalization and Adaptation | ||||||
[168] | Personalized Generation | SD | 3-5 concept images + text prompt | Concept-specific generation with few shots | Context-appearance entanglement issues | Class-specific prior preservation loss |
[68] | Personalized Generation | SD v1.4 | 3-5 concept images | Single word embedding for concept representation | Lengthy optimization (2h per concept) | 5000 optimization steps with reconstruction loss |
[157] | Personalized Generation | SD 1.5 | 5-50 concept images + text prompt | Orthogonal transformation preserving hyperspherical energy | Color and detail distortions | Orthogonal weight transformation with energy preservation |
[138] | Editing | SD v1 | Image + source caption | Null-text optimization without model tuning | Requires editable parts in source caption | - |
[264] | Style Transfer | SD | Content image + style image | Single-image style learning via attention | Long training time (20min/image) | - |
[227] | Style Transfer | SD v1.4 | Image + text prompt | Text embedding optimization for attribute editing | Struggles with fine-grained edits | No fine-tuning, only embedding optimization |
[13] | Classification | SD 1.5 | Class labels or images | Zero-shot data generation for robustness | Limited to ImageNet-like datasets | - |
[262] | Classification | SD 1.4, DALL-E2 | Small-scale seed dataset | Guided imagination for class-consistent generation | Generated samples less informative than real ones | Domain adaptation fine-tuning |
Application Specific Augmentation | ||||||
[177] | Brain Lesion Detection | - | Multi-channel MRI scans | Counterfactual healthy image generation | Limited to brain tumor scenarios | Healthy/unhealthy classification training |
[155] | Brain MRI Generation | - | Age, sex, brain metrics | Anatomically-controlled brain MRI synthesis | Limited to training distribution range | Conditional training on demographic variables |
[61] | Object Detection | SD | Visual prior + text prompt | Category-calibrated CLIP filtering | Performance drops with excessive synthetic data | - |
From the observation of Table 3, several key insights can be derived. In particular, SD is shown to be the most widely used FDM across multiple tasks, mainly due to its open-source characteristic. Additionally, an important trade-off between processing speed and output quality is observed, where methods requiring optimization or fine-tuning typically produce better results, but at the cost of longer processing times (often ranging from minutes to hours per image). Another notable trend is the evolution from the usage of single- to multi-condition mechanisms, where methods increasingly support various input modalities (i.e., text, images, masks, sketches) for more precise control; however, this increased flexibility is often accompanied with higher computational requirements and more complex training procedures. Investigating category-specific patterns, semantic manipulation methods typically focus on architectural innovations and generally require less fine-tuning, often implementing zero- or few-shot techniques; these methods also emphasize user control through various input modalities. On the other hand, personalization and adaptation methods demonstrate a clear trend towards optimization-heavy approaches, exhibiting notably longer processing times (often at the scale of hours) and more complex fine-tuning requirements, in order to preserve identity and style characteristics. Moreover, application-specific methods are characterized by their specialized nature, requiring domain-specific training data and often implementing custom architectures or loss functions tailored to their selected application fields; these methods typically prioritize task-specific performance metrics over general-purpose applicability.